solve by completing the square


3x^2 +5x=2


answer: -2,1/3 can anyone give me full break down on why the answer is -2,1/3

Answer by scott8148(6628)

dividing by 3 ___ x^2 + 5/3 x = 2/3

adding half of the x-term coefficient, squared ___ x^2 + 5/3 x + (5/6)^2 = 49/36

taking square root ___ x + 5/6 = ± 7/6

x + 5/6 = 7/6 ___ x = 1/3

x + 5/6 = -7/6 ___ x = -2